# CoreMark: Toward Robust and Universal Text Watermarking Technique

### v1

@@ Abstract
Text watermarking schemes have gained considerable attention in recent years, yet still face critical challenges in achieving simultaneous robustness, generalizability, and imperceptibility. This paper introduces a new embedding paradigm, termed core, which comprises several consecutively aligned black pixel segments. Its key innovation lies in its inherent noise resistance during transmission and broad applicability across languages and fonts. Based on the core, we present a text watermarking framework named CoreMark. Specifically, CoreMark first dynamically extracts cores from characters. Then, the characters with stronger robustness are selected according to the lengths of cores. By modifying the thickness of the core, the hidden data is embedded into the selected characters without causing significant visual distortions. Moreover, a general plug-and-play embedding strength modulator is proposed, which can adaptively enhance the robustness for small font sizes by adjusting the embedding strength according to the font size. Experimental evaluation indicates that CoreMark demonstrates outstanding generalizability across multiple languages and fonts. Compared to existing methods, CoreMark achieves significant improvements in resisting screenshot, print-scan, and print-camera attacks, while maintaining satisfactory imperceptibility.
@@ I Introduction
Based on the core concept, we propose an image-based text watermarking method (dubbed ‘CoreMark’), which consists of two main stages: watermark embedding and extraction. Specifically, during the watermark embedding stage, information is embedded into characters by modifying pixels to adjust the thickness of core. Unlike edge features that rely on contour details, thickness is a structural attribute that remains preserved after binarization. In the extraction stage, an adaptive threshold is first computed, and the embedded information is then recovered by comparing the width of each character’s core against this threshold. In particular, we take the human perceptual factor into account by referring to the neighboring pixels of core to ensure the structural consistency of characters before and after embedding, thereby guaranteeing invisibility. Besides the improvements mentioned above, we observe that robustness declines significantly as font size decreases [12, 8]. To enhance robustness, traditional methods typically employ a fixed repetition strategy for embedding. However, this approach may offer little to no additional benefit for larger characters that are already sufficiently robust, and can even compromise imperceptibility. To address this limitation, we propose a general plug-and-play embedding strength (ES) modulator that adaptively increases the embedding strength as the font size decreases.
@@
Image-based Approach. Image-based approaches [7, 16, 17, 18, 19, 9, 8, 20] handle characters as image data, hiding information by manipulating individual pixels to transform the statistical features of the characters. Wu and Liu [7] firstly established a set of flipping rules and proposed a method for embedding watermarks by flipping pixels, laying an important theoretical foundation for subsequent research[16, 17, 18, 19]. However, the authors [7] state that these approaches are only effective under high-resolution printing and scanning conditions, such as the 9600$\times$2400 dpi printer used in [17]. Yang et al. [9] initially embedded data based on altering the relative center of gravity between paired characters, which shows strong language independence. Despite various efforts, a clear trade-off remains between robustness and invisibility in existing binary image watermarking methods. Methods that improve robustness typically incur visible distortions, whereas those designed for better imperceptibility often fail to withstand common degradations such as printing and scanning. This imbalance highlights the difficulty of simultaneously achieving both properties, especially in real-world degradation scenarios.
@@ III-A Preliminaries
Figure 3: Robustness of methods [7, 8, 9] without attack.
Representative Methods in our Evaluation. We evaluated three image-based text watermarking methods: Wu04 [7], Tan19 [8], and Yang23 [10]. Wu04 embeds watermarks by pixel flipping to ensure the black pixel count in $s\times s$ blocks equals $2kQ$ for bit ‘0’ and $(2k+1)Q$ for bit ‘1’. Tan19 modifies character boundary pixels to achieve the black pixel count $\gamma$. Specifically, the rounded value of $\gamma/(K\times\Theta)$, where $\Theta$ denotes the average black pixels per character, must be odd for bit ‘0’ and even for bit ‘1’. Yang23 manipulates relative centroid positions of adjacent characters: a bit ‘0’ is embedded when the left character’s centroid exceeds the right one by a threshold $R$, and bit ‘1’ otherwise. When embedding conditions are not met, the pixels of both characters are adjusted by $r_{\text{diff}}/2$ to satisfy requirements. Here, the parameters $Q$, $K$, and $R$ are quantization steps that determine the embedding strength, with higher values improving robustness while degrading imperceptibility.
@@ III-B Challenge 1: Invisibility-Robustness Tradeoff
Settings. To quantify this trade-off, we evaluate three methods across 50 English documents rendered in Arial font, assessing imperceptibility and robustness against screenshot attacks under varying quantization sizes (denoted as Step 1 through Step 5). Specifically, Wu04 employs steps from 100 to 500 (interval: 100), Tan19 uses 0.05 to 0.25 (interval: 0.05), and Yang23 utilizes 6 to 14 (interval: 2). We measure robustness via accuracy ($ACC$) defined in Eq. 1:
@@
Results. As Figure 3 illustrates, Wu04’s $ACC$ consistently hovers around 50% regardless of the step size, equivalent to random guessing. This results from screenshot-induced resizing effects disrupting encoder-decoder synchronization due to unknown resizing factors. For Tan19, increasing the quantization step significantly improves robustness, achieving over 90% accuracy at Step 5. However, this comes at the cost of severe imperceptibility degradation, with PSNR dropping below the commonly accepted threshold of 30 dB [38, 39]. Yang23 consistently produces PSNR values below 30 dB due to extensive pixel modifications required toto meet embedding thresholds. Meanwhile, its accuracy remains below 90%, primarily due to embedding instability (discussed in Section III-C) and susceptibility to noise. These results reveal two fundamental design principles for watermarking methods: (1) robustness is inherently linked to a method’s ability to withstand scaling distortions commonly encountered in real-world transmission scenarios, such as screen capture and image resizing. Consequently, resistance to such transformations must be explicitly incorporated into method design; and (2) for image-based methods, an effective approach to mitigating the robustness-imperceptibility trade-off is to develop novel embedding paradigms that enable information embedding through modification of only a small number of pixels. This approach ensures that even with increased embedding strength, visual degradation remains acceptable.
@@ III-C Challenge 2: Embedding Stability
Settings. We also use Wu04, Tan19, and Yang23 in 50 documents for our analysis. The quantization step settings are identical to those described in Section III-B. To evaluate embedding stability, we first embed the same watermark sequences at varying step sizes, then extract them without attacks, and compute $ACC$, where a lower value indicates greater embedding instability.
Results. As shown in Figure 3, Tan19 shows no signs of embedding instability. Conversely, Wu04’s instability increases as the quantization step size increases. This is due to the fact that a larger quantization step size requires more pixel modifications. When the number of modifiable pixels is limited, some characters cannot meet the embedding conditions, resulting in embedding failure. Embedding failures in Yang23 occur only for the characters ‘l’ and ‘i’. These symmetric characters maintain centered centroids even after modifications, limiting the centroid difference to $r_{\text{diff}}/2$ relative to adjacent characters. As a result, they fail to satisfy the embedding criteria, leading to embedding instability. These observations suggest two key insights: (1) ensuring all characters are suitable for embedding is challenging due to diverse character shapes and pixel distributions; and (2) developing methods that enable flexible character selection for embedding is more practical.
@@ III-D Challenge 3: Resistance to Binarization
Results. As shown in Figure 4, while AutoStegaFont achieves nearly perfect accuracy under clean and screenshot conditions without binarization, its accuracy drops by nearly 20% when binarization is applied. Similar degradation is observed under print–scan and print–camera scenarios, where binarization reduces $ACC$ by over 10%. These results highlight the limitations of methods that relying heavily on edge features, which are susceptible to degradation induced by binarization.
@@ IV The Proposed Method
To address these challenges, we first introduce an embedding paradigm, termed core, based on which we introduce CoreMark, a novel watermarking method. As illustrated in Figure 5, CoreMark operates in two distinct stages: watermark embedding and watermark extraction.
@@ IV-A Watermark Embedding
Component 1: core and its Extraction Model. A core comprises a series of consecutive black pixels, denoted as $S_{\text{core}}=\left\{S_{i}\mid i=1,\dots,n\right\}$, where $n$ represents the total number of black pixel sequences. Each core is determined by three fundamental parameters: direction $\mathcal{D}$, starting point $\mathcal{P}=\left\{\left(x_{i},y_{i}\right)\mid i=1,\dots,n\right\}$, and length =$\mathcal{L}=\left\{l_{i}\mid i=1,\dots,n\right\}$. Specifically, when direction $\mathcal{D}$ is horizontal, each $S_{i}$ is extracted from an image row; when $\mathcal{D}$ is vertical, each $S_{i}$ is extracted from a column. The starting point $\mathcal{P}$ indicates the initial pixel in each $S_{i}$, corresponding to the leftmost pixel when $\mathcal{D}$ is horizontal and the topmost pixel when $\mathcal{D}$ is vertical. The length $\mathcal{L}$ specifies the number of black pixels in each sequence. In the following, we provide a detailed explanation of the calculation methods for these three parameters, separately.
@@
(1) Determine the direction $\mathcal{D}$. We determine $\mathcal{D}$ by comparing the average length of consecutive black pixels across different directions. Taking the horizontal direction as an example, let the dimensions of the character image be (M $\times$ N). For each row $R_{m},m=1,...,M$, we apply Run Length Coding (RLC) [40] to represent $R_{m}$ as a series of data pairs $\zeta_{m}(\mu_{j}^{m},\rho_{j}^{m})$, where each $\zeta_{m}(\mu_{j}^{m},\rho_{j}^{m})$ signifies a sequence of consecutive pixels with identical values. Here, $\mu_{j}^{m}\in(0,\rm N]$ is the repetition count and $\rho_{j}^{m}\in\left\{0,1\right\}$ denotes the pixel value being repeated (where ”0” represents black pixels and ”1” represents white pixels). For instance, given the following pixel sequence for $R_{i}$:
$11,1,1,1,\underbrace{0,\cdots,0}_{36\enspace pixels},1,1,$ 1 (2)
the resulting data pairs are $\zeta_{m}(4,1)$, $\zeta_{m}(36,0)$, $\zeta_{m}(3,1)$. For each $R_{m}$, we identify the maximum value of $\mu_{j}^{m}$ across all pairs, denote as $\mu_{\text{max}}^{m}$, with the corresponding $\rho_{j}^{m}$ denoted as $\rho_{\text{max}}^{m}$. After determining these values for all rows, we compute the average length of consecutive black pixels in the horizontal direction by dividing $U_{h}$ by $K_{h}$, where $U_{h}$ and $K_{h}$ are calculated as follows.
$U_{h}=\sum_{r=1}^{M}{\mu_{\text{max}}^{m}\times{\rm sign_{v}}\left(\rho_{\text{max}}^{m}\right)}.$ (3)
$\rm sign_{v}(\cdot)$ is the sign function defined by
@@
$K_{h}$ represents the frequency of occurrences where $\rho_{\text{max}}^{m}=0$ across all rows. We apply identical procedures to derive $U_{v}$ and $K_{v}$ for the vertical direction. If $\frac{U_{h}}{K_{h}}$ exceeds $\frac{U_{v}}{K_{v}}$, the direction is horizontal, and $S_{\text{core}}$ is calculated using row pixels; otherwise, $S_{\text{core}}$ is calculated using column pixels. Figure 6 shows the $U_{h}$ calculation process in the horizontal direction.
(2) Determine the starting coordinates $\mathcal{P}$ and lengths $\mathcal{L}$. In general, multiple consecutive black pixel sequences exist along $\mathcal{D}$; however, not all consecutive black pixels along $\mathcal{D}$ qualify as components of $S_{\text{core}}$. Below, we describe the selection process for suitable consecutive black pixel segments in the horizontal direction to construct $S_{\text{core}}$. We assign the pixel sequences $\zeta_{m}(\mu_{\text{max}}^{m},\rho_{\text{max}}^{m})$ in $\mathcal{D}$, where $\rho_{\text{max}}^{m}=0$, as ‘candidate vectors’, denoted as $V_{\text{cand}}$. The corresponding $m$ values of all ‘candidate vectors’ are in the array $\varPhi$. The starting coordinates and lengths of $V_{\text{cand}}$ are represented as $C_{\text{start}}=\left\{\left(x_{t},y_{t}\right)|\ t\in\varPhi\right\}$ and $C_{\text{len}}=\left\{\mu_{t}^{\text{max}}|\ t\in\varPhi\right\}$. The coordinates $(x_{t},y_{t})$ can be directly derived from $\zeta_{m}(\mu_{\text{max}}^{m},\rho_{\text{max}}^{m})$, where $x_{t}=m$ represents the row index of the sequence $\mu_{\text{max}}^{m}$. $y_{t}$ can be computed as follows:
$j_{\text{max}}=\mathop{\arg\max}\limits_{j}\mu_{\text{max}}^{m};\quad y_{t}=\sum\nolimits_{t=1}^{j_{\text{max}}-1}{\mu_{t}^{m}}+1.$ (5)
Next, we cluster all ‘candidate vectors’ based on their $C_{\text{start}}$ and $C_{\text{len}}$, subsequently selecting one cluster as the final $S_{\text{core}}$.
Algorithm 1 Method of assigning “candidate vectors”.
0:  $C_{\text{start}}$, $C_{\text{len}}$
0:  $\mathcal{I}$
1:  Initiate $C_{1,...,k}^{\text{num}}$, $C_{1,...,k}^{\text{ave}}$
2:  ///// / $C_{1,...,k}^{\text{num}}$ is the number of “candidate vectors” in each cluster; $C_{1,...,k}^{\text{ave}}$ is the average length of each cluster.
@@
4:  $\hat{y}\leftarrow y_{1}$, $\hat{x}\leftarrow x_{1}$, $\mu_{\text{max}},\mu_{\text{joint}}\leftarrow\mu_{x_{1}}^{\text{max}}$, $n\leftarrow 1$, $k\leftarrow 0$
5:  for each $i\in[2,P_{r}]$ do
6:     ///// / Verify compliance with the two rules.
7:     $b_{1}\leftarrow x_{i}-\hat{x}==1$, $b_{2}\leftarrow{\rm abs}(y_{i}-\hat{y})\leq T_{c}$
8:     $b_{3}\leftarrow{\rm abs}(y_{i}+\mu_{x_{i}}^{\text{max}}-\hat{y}-\mu_{\text{max}})\leq T_{c}$
9:     if $b_{1}\&b_{2}\&b_{3}$ then
10:        ///// / If it matches, join the current cluster.
11:        $\hat{x}\leftarrow x_{i}$ , $\hat{y}\leftarrow y_{i}$, $\mu_{\text{max}}\leftarrow\mu_{x_{i}}^{\text{max}}$, $\mu_{\text{joint}}\leftarrow\mu_{\text{joint}}+\mu_{x_{i}}^{\text{{max}}}$, $n\leftarrow n+1$
12:     else
13:        ///// / If it doesn’t match, start the next cluster.
14:        $C_{k}^{\text{num}}\leftarrow n$, $C_{k}^{\text{ave}}\leftarrow\frac{\mu_{\text{joint}}}{n}$, $k\leftarrow k+1$, $n\leftarrow 1$, $\mu_{\text{max}},\mu_{\text{joint}}\leftarrow\mu_{x_{i}}^{\text{max}}$, $\hat{x}\leftarrow x_{i}$ , $\hat{y}\leftarrow y_{i}$
15:     end if
16:  end for
17:  $k\leftarrow k+1$, $C_{k}^{\text{num}}\leftarrow n$, $C_{k}^{\text{ave}}\leftarrow\frac{\mu_{\text{joint}}}{n}$, $k_{\text{max}}\leftarrow\mathop{\arg\max}_{k}(C_{k}^{\text{ave}})$
18:  $\beta^{\text{start}}\leftarrow\sum_{t=1}^{k_{\text{max}}}{C_{t}^{\text{num}}}$, $\beta^{\text{end}}\leftarrow\sum_{t=1}^{k_{\text{max}}}{C_{t}^{\text{num}}}+C_{k_{\text{max}}}^{\text{num}}$
19:  Put $y_{\beta^{\text{start}}},...,y_{\beta^{\text{end}}}$ in list $\mathcal{I}$
20:  return  $\mathcal{I}$
The clustering rules for ‘candidate vectors’ are as follows: (1) rows of $V_{\text{cand}}$s within a cluster must be adjacent; (2) the distance between $y_{t}$ values of each $V_{\text{cand}}$ in a cluster must be less than $T_{c}$; (3) the distance between $y_{t}+\mu_{x_{t}}^{\text{max}}$ values must also be less than $T_{c}$. After applying these rules and generating multiple clusters, we select the cluster with the highest $C_{\text{max}}^{\text{ave}}$ as the final $S_{\text{core}}$, where $C_{\text{max}}^{\text{ave}}$ represents the average value of $\mu_{x_{t}}^{\text{max}}$ within each cluster. We use $\mathcal{I}$ to store the indices of $R_{i}$s where the ’candidate vectors’ of the final selected cluster are located. With $\mathcal{I}$, we can readily extract the starting coordinates $\mathcal{P}$ and lengths $\mathcal{L}$ of $S_{\text{core}}$ from $C_{\text{start}}$ and $C_{\text{len}}$, thereby obtaining the three essential parameters for locating $S_{\text{core}}$, namely direction, starting coordinates, and length. The same methodology applies when the direction is vertical, with the corresponding $S_{\text{core}}$ determined through identical steps. The complete procedure is depicted in Algorithm 1.
We now explain the robustness of $S_{\text{core}}$, which arises from two key factors. 1) We exclusively focus on $S_{\text{core}}$, ensuring that unconnected noisy pixels exert no influence on its extraction. 2) We employ RLC to represent each $R_{i}$, ensuring that only sequences of consecutive pixels with identical values are considered valid candidates $V_{\text{cand}}$s. Since noisy pixels typically appear as isolated and randomly distributed points, they are inherently excluded from $V_{\text{cand}}$.

### v2

@@ Abstract
Text watermarking schemes have gained considerable attention in recent years, yet still face critical challenges in achieving simultaneous robustness, generalizability, and imperceptibility. This paper introduces a new embedding paradigm, termed core, which comprises several consecutively aligned black pixel segments. Its key innovation lies in its inherent noise resistance during transmission and broad applicability across languages and fonts. Based on the core, we present a text watermarking framework named CoreMark. Specifically, CoreMark first dynamically extracts cores from characters. Then, the characters with stronger robustness are selected according to the lengths of cores. By modifying the size of the core, the hidden data is embedded into the selected characters without causing significant visual distortions. Moreover, a general plug-and-play embedding strength modulator is proposed, which can adaptively enhance the robustness for small font sizes by adjusting the embedding strength according to the font size. Experimental evaluation indicates that CoreMark demonstrates outstanding generalizability across multiple languages and fonts. Compared to existing methods, CoreMark achieves significant improvements in resisting screenshot, print-scan, and print-camera attacks, while maintaining satisfactory imperceptibility.
@@ I Introduction
Based on the core concept, we propose an image-based text watermarking method (dubbed ‘CoreMark’), which consists of two main stages: watermark embedding and extraction. Specifically, during the watermark embedding stage, information is embedded into characters by modifying pixels to adjust the size of core. Unlike edge features that rely on contour details, the size of core is a structural attribute that remains preserved after binarization. In the extraction stage, an adaptive threshold is first computed, and the embedded information is then recovered by comparing the width of each character’s core against this threshold. In particular, we take the human perceptual factor into account by referring to the neighboring pixels of core to ensure the structural consistency of characters before and after embedding, thereby guaranteeing invisibility. Besides the improvements mentioned above, we observe that robustness declines significantly as font size decreases [12, 8]. To enhance robustness, traditional methods typically employ a fixed repetition strategy for embedding. However, this approach may offer little to no additional benefit for larger characters that are already sufficiently robust, and can even compromise imperceptibility. To address this limitation, we propose a general plug-and-play embedding strength (ES) modulator that adaptively increases the embedding strength as the font size decreases.
@@
[…]
@@ III-A Preliminaries
Representative Methods in our Evaluation. We evaluated three image-based text watermarking methods: Wu04 [7], Tan19 [8], and Yang23 [10]. Wu04 embeds watermarks by pixel flipping within fixed-size blocks to make black pixel counts in blocks even or odd multiples of quantization step $Q$ for bits ‘0’ and ‘1’, respectively. Tan19 modifies character boundary pixels to control the total black pixel count within each character, embedding bits based on whether the rounded ratio of black pixels to average pixels per character (scaled by $K$) is odd or even. Yang23 manipulates relative centroid positions of adjacent characters, embedding bit 0 when the left centroid exceeds the right by threshold $R$, otherwise bit 1. The parameters $Q$, $K$, and $R$ serve as quantization steps that control embedding strength. Higher values improve robustness against attacks but reduce visual imperceptibility.
@@ III-B Challenge 1: Invisibility-Robustness Tradeoff
Settings. To quantify this trade-off, we evaluate three methods across 50 English documents rendered in Arial font. We assess imperceptibility and robustness against screenshot attacks under five different quantization steps. Specifically, Wu04 employs $Q$ values from 100 to 500 (interval: 100), Tan19 uses $K$ values from 0.05 to 0.25 (interval: 0.05), and Yang23 utilizes $R$ values from 6 to 14 (interval: 2). Robustness is measured via accuracy (ACC) defined as:
@@
Results. As Figure 3 illustrates, Wu04’s $ACC$ consistently hovers around 50% regardless of the step size, equivalent to random guessing. This results from screenshot-induced resizing effects disrupting encoder-decoder synchronization due to unknown resizing factors. For Tan19, increasing the quantization step significantly improves robustness, achieving over 90% accuracy at Step 5. However, this comes at the cost of severe imperceptibility degradation, with PSNR dropping below the commonly accepted threshold of 30 dB [38, 39]. Yang23 consistently produces PSNR values below 30 dB due to extensive pixel modifications required to meet embedding thresholds. Meanwhile, its accuracy remains below 90%, primarily due to embedding instability (discussed in Section III-C) and susceptibility to noise. These results reveal two fundamental design principles for watermarking methods: (1) robustness is inherently linked to a method’s ability to withstand scaling distortions commonly encountered in real-world transmission scenarios, such as screen capture and image resizing. Consequently, resistance to such transformations must be explicitly incorporated into method design; and (2) for image-based methods, an effective approach to mitigating the robustness-imperceptibility trade-off is to develop novel embedding paradigms that enable information embedding through modification of only a small number of pixels. This approach ensures that even with increased embedding strength, visual degradation remains acceptable.
@@ III-C Challenge 2: Embedding Stability
Settings. We evaluate embedding instability using the same three methods and 50 documents from Section III-B. We embed identical watermark sequences at different parameter values, extract them without attacks, and measure $ACC$. Lower $ACC$ values indicate greater embedding instability.
Results. As shown in Figure 3, Tan19 exhibits stable embedding across all parameter values. Conversely, Wu04’s instability increases as the quantization step size increases. This is due to the fact that a larger quantization step size requires more pixel modifications. When the number of modifiable pixels is limited, some characters cannot meet the embedding conditions, resulting in embedding failure. Yang23 experiences embedding failures primarily with narrow characters like ‘l’ and ‘i’. These symmetric characters maintain centered centroids even after modifications, limiting the centroid difference to the embedding threshold relative to adjacent characters. As a result, they fail to satisfy the embedding criteria, leading to embedding instability. These observations suggest two key insights: (1) ensuring all characters are suitable for embedding is challenging due to diverse character shapes and pixel distributions; and (2) developing methods that enable flexible character selection for embedding is more practical.
@@ III-D Challenge 3: Resistance to Binarization
Results. As shown in Figure 4, AutoStegaFont achieves nearly perfect accuracy under clean and screenshot conditions without binarization. However, its accuracy drops by nearly 20% when binarization is applied. Similar degradation is observed under print–scan and print–camera scenarios, where binarization reduces $ACC$ by over 10%. These results highlight the vulnerability of methods that rely heavily on edge features, which are susceptible to degradation induced by binarization.
@@
TABLE I: The Notations of Explanation.
Notation Explanation
$\zeta_{m}(\mu^{j}_{m},\rho^{j}_{m})$ A consecutive sequence of pixels in row $m$ with
length $\mu^{j}_{m}$ and pixel value $\rho^{j}_{m}$.
$\zeta_{m}(\mu_{m}^{j_{\max}},\rho_{m}^{j_{\max}})$ The longest consecutive pixel sequence in row $m$ with
length $\mu_{m}^{j_{\max}}$ and pixel value $\rho_{m}^{j_{\max}}$.
$S_{\text{core}}$ core of each character
$\mathcal{I}$ The indices of rows belonging to $S_{\text{core}}$.
$\mathcal{D},\mathcal{P},\mathcal{L}$ Direction, starting point, and length of $S_{\text{core}}$, respectively.
$V_{\text{cand}}$ Row indices of the candidate vectors.
$\varPhi$ The set of row indices of candidate vectors.
$C_{\text{start}}$, $C_{\text{len}}$ The starting coordinates and lengths of candidate vectors.
$C$ Set of all characters.
$n_{\text{core}}$, $\tilde{n}_{\text{core}}$ The size of core of each $c_{i}$ before and after embedding.
$T_{\Delta}$ The embedding threshold.
$\beta$ The watermark embedding strength of CoreMark.
@@ IV The Proposed Method
To address the above challenges, we first introduce an embedding paradigm, termed core, based on which we introduce CoreMark, a novel watermarking method. As illustrated in Figure 5, CoreMark operates in two distinct stages: watermark embedding and watermark extraction. The notations within this paper are shown in Table I.
@@ IV-A Watermark Embedding
Component 1: core and its Extraction Model. A core, denoted as $S_{\text{core}}$, comprises a series of consecutive black pixels. Each core is determined by three fundamental parameters: direction $\mathcal{D}\in\{\text{horizontal, vertical}\}$, starting point $\mathcal{P}=\{(x_{i},y_{i})\}_{i=1}^{n_{\text{core}}}$, and length $\mathcal{L}=\{\ell_{i}\}_{i=1}^{n_{\text{core}}}$, where $n_{\text{core}}$ is the total number of such sequences. Specifically, $S_{\text{core}}$ is extracted from rows if $\mathcal{D}$ is horizontal, or from columns if $\mathcal{D}$ is vertical. The starting point $\mathcal{P}$ indicates the initial pixel of each sequence, corresponding to the leftmost pixel when $\mathcal{D}$ is horizontal and the topmost pixel when $\mathcal{D}$ is vertical. The length $\mathcal{L}$ specifies the number of black pixels in each sequence. The calculation methods for these three parameters ($\mathcal{D},\mathcal{P},\mathcal{L}$) are explained in detail in the following subsections.
@@
(1) Determine the direction $\mathcal{D}$. Let the character image have dimensions $M\times N$. Taking the horizontal direction as an example, let $R_{m}$ denote the $m$-th row of the image, $m=1,\dots,M$. For each row $R_{m}$, we apply Run Length Coding (RLC) [40] to represent $R_{m}$ as a series of data pairs $\zeta_{m}(\mu^{j}_{m},\rho^{j}_{m}),j=1,\dots,J_{m}$, where $J_{m}\in\{1,\dots,N\}$. Here, $\mu^{j}_{m}\in\{1,\dots,N\}$ is the length of a consecutive pixel sequence, and $\rho^{j}_{m}\in\{0,1\}$ indicates the pixel value (0 for black, $1$ for white). Let $\mu_{m}^{j_{\max}}$ denote the length of the longest consecutive pixel sequence in row $m$, and $\rho_{m}^{j_{\max}}$ its pixel value. The direction $\mathcal{D}$ is determined as follows:
$\mathcal{D}=\begin{cases}\text{horizontal},&\text{if }\dfrac{U_{h}}{K_{h}}>\dfrac{U_{v}}{K_{v}},\\[2.84526pt] \text{vertical},&\text{otherwise.}\end{cases}$ (2)
where $U_{h}$ is calculated by
$U_{h}=\sum_{m=1}^{M}{\mu_{m}^{j_{\max}}\times\mathbf{1}_{\{0\}}\left(\rho_{m}^{j_{\max}}\right)}.$ (3)
The indicator function $\mathbf{1}_{\{0\}}$ is defined as:
$\mathbf{1}_{\{0\}}(x)=\begin{cases}1,&x=0,\\ 0,&x=1.\end{cases}$ (4)
$K_{h}=\sum_{m=1}^{M}\mathbf{1}_{\{0\}}(\rho_{m}^{j_{\max}})$ represents the number of rows with $\rho_{m}^{j_{\max}}=0$. Analogous quantities $U_{v}$ and $K_{v}$ are computed along the vertical direction. Figure 6 illustrates the calculation process of $U_{h}$ in the horizontal direction.
(2) Determine the starting coordinates $\mathcal{P}$ and lengths $\mathcal{L}$. In general, multiple consecutive black pixel sequences exist along $\mathcal{D}$; however, not all of them qualify as elements of $S_{\text{core}}$. We illustrate the selection process in the horizontal case. We first identify candidate vectors from the RLC output. For each row, the longest black pixel run $\zeta_{m}(\mu_{m}^{j_{\max}},\rho_{m}^{j_{\max}})$ with $\rho_{m}^{j_{\max}}=0$ is designated as candidate vector. The set of candidate vectors is denoted by $V_{\text{cand}}$, with their row indices collected in $\varPhi$. The starting coordinates and lengths of $V_{\text{cand}}$s are defined as $C_{\text{start}}=\{(x_{t},y_{t})\}_{t=1}^{\text{len}(\varPhi)}$ and $C_{\text{len}}=\{\mu_{t}^{j_{\max}}\}_{t=1}^{\text{len}(\varPhi)}$. Here, $x_{t}=m$ is the row index, and its corresponding $y_{t}$ gives the starting column of the longest run, which is computed as follows.
$y_{t}=\sum\nolimits_{i=1}^{{j_{\max}-1}}{\mu^{i}_{m}}+1.$ (5)
Next, the set of candidate vector is clustered according to spatial continuity.
Algorithm 1 Method of assigning candidate vectors.
@@
[…]
1: Initiate $C_{\text{num}}=\{C_{i}^{\text{num}}\}_{i=1}^{k}$, $C_{\text{ave}}=\{C_{i}^{\text{ave}}\}_{i=1}^{k},k\in\mathbb{Z}^{+}$
2://// $C_{i}^{\text{num}}$ is the number of candidate vectors in each cluster; $C_{i}^{\text{ave}}$ is the average length of each cluster.
@@
[…]
5:for each $i\in[2,\text{len}(\varPhi)]$ do
6:  //// Verify compliance with three rules.
7:  $b_{1}\leftarrow x_{i}-\hat{x}==1$
8:  $b_{2}\leftarrow{\rm abs}(y_{i}-\hat{y})\leq T_{c}$
9:  $b_{3}\leftarrow{\rm abs}(y_{i}+\mu_{x_{i}}^{j_{\max}}-\hat{y}-\mu^{j_{\max}}_{\hat{x}})\leq T_{c}$
10:  if $b_{1}\&b_{2}\&b_{3}$ then
11:   //// If it matches, join the current cluster.
12:   $\hat{x}\leftarrow x_{i}$ , $\hat{y}\leftarrow y_{i}$, $\mu_{\text{joint}}\leftarrow\mu_{\text{joint}}+\mu_{x_{i}}^{j_{\max}}$, $n\leftarrow n+1$
13:  else
14:   //// If it doesn’t match, start the next cluster.
15:   $C_{k}^{\text{num}}\leftarrow n$, $C_{k}^{\text{ave}}\leftarrow\frac{\mu_{\text{joint}}}{n}$, $\mu_{\text{joint}}\leftarrow\mu_{x_{i}}^{j_{\max}}$, $\hat{x}\leftarrow x_{i}$ , $\hat{y}\leftarrow y_{i}$, $k\leftarrow k+1$, $n\leftarrow 1$
16:  end if
17:end for
[…]
20:return  $\mathcal{I}=\{x_{s},x_{s}+1,\dots,x_{e}\}$
The clustering is constrained by three rules: (1) the rows of $V_{\text{cand}}$s within a cluster must be adjacent; (2) the distance between their $y_{t}$ values must be less than $T_{c}$; (3) the distance between their $y_{t}+\mu_{x_{t}}^{j_{\max}}$ values must also be less than $T_{c}$. By applying the above procedure, multiple clusters can be obtained. The length of each cluster is subsequently computed, and the cluster with the maximum length is selected as the final $S_{\text{core}}$. The set $\mathcal{I}=\{\mathcal{I}_{i}\}_{i=1}^{n_{\text{core}}},\mathcal{I}_{i}\in\{1,\dots,M\}$ records the indices of rows belonging to $S_{\text{core}}$, from which the starting coordinates $\mathcal{P}$ and lengths $\mathcal{L}$ are directly extracted from $C_{\text{start}}$ and $C_{\text{len}}$. The complete procedure is depicted in Algorithm 1.
The robustness of $S_{\text{core}}$ is attributed to two factors. 1) We exclusively focus on $S_{\text{core}}$, ensuring that unconnected noisy pixels exert no influence on its extraction. 2) The use of RLC guarantees that only consecutive runs of identical pixel values are considered valid candidates $V_{\text{cand}}$s. Since noise typically manifests as isolated pixels, such disturbances are inherently eliminated during the construction of $V_{\text{cand}}$.